solving quadratic equations by completing the square

HELP PLEASE
im so confusedddddd -3 doesnt add up to 1

help???

x^2+x-3=0

thank you!! will give beautiful review

Answer by mananth(16946)

x^2+x-3=0
x^2+x=3
x^2+x+1/4 = 3+1/4
(x+1/2)^2 = 13/4
take the square root
x+1/2 = +-sqrt(13)/2
x=sqrt(13)/2 -1/2
OR
x=-sqrt(13)/2-1/2